OpenStudy (anonymous):

Which ordered pair in the form (a,b) is a solution of this equation? 3a-4b=21 A) (7,0) B) (-2,-3) C) (0,7) D) (-3,-2)

OpenStudy (anonymous):

@Peach-es1234

OpenStudy (alexandervonhumboldt2):

subtitute the points: 3*7-4*0=21 check if this is correct 3*(-2)-4*(-3)=21 check is this is correct 3*0-4*7=21 check is this is correct 3*(-3)-4*(-2)=21 check is this is correct

OpenStudy (vickyisthesmartone123):

Well fit in each number and it should give you the answer

OpenStudy (anonymous):

C - (7,0) because 3x7= 21 add 4x0 = 0 which then equals 21 :)

OpenStudy (vickyisthesmartone123):

Nope it is the second one

OpenStudy (anonymous):

ummm....?? i dont kow who is right? @Vickyisthesmartone123 can you explain?

OpenStudy (alexandervonhumboldt2):

3*7-4*0=21 3*7-0=21 21=21

OpenStudy (anonymous):

@AlexandervonHumboldt2 yea i was kinda leaning toward that one since the others didnt really make sense to me thx

OpenStudy (anonymous):

It is a! :)

OpenStudy (vickyisthesmartone123):

Yep sorry it is A

OpenStudy (anonymous):

@Peach-es1234 thx for your help. another?

OpenStudy (anonymous):

Ya I put c instead of a. sorry bout the confusion.

OpenStudy (anonymous):

Ask Away! :)

OpenStudy (vickyisthesmartone123):

MEDAL ME OR WHOEVER HELPED YOU THE MOST KID!